OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, –1). Im having trouble getting this can you explain this to me with a formula to use

OpenStudy (anonymous):

y = -1/2x + 6

OpenStudy (anonymous):

y = x − 3 y = x + 17 y = –3x − 3 y = –3x + 17 these are my only options to answer its really weird

OpenStudy (anonymous):

oh sorry, then the answer is A.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

welcome! @Giovanni